### v1

# TimeCopilot

Abstract
We introduce TimeCopilot, the first open-source agentic framework for forecasting that combines multiple Time Series Foundation Models (TSFMs) with Large Language Models (LLMs) through a single unified API. TimeCopilot automates the forecasting pipeline: feature analysis, model selection, cross-validation, and forecast generation, while providing natural language explanations and supporting direct queries about the future. The framework is LLM-agnostic, compatible with both commercial and open-source models, and supports ensembles across diverse forecasting families. Results on the large-scale GIFT-Eval benchmark show that TimeCopilot achieves state-of-the-art probabilistic forecasting performance at low cost. Our framework provides a practical foundation for reproducible, explainable, and accessible agentic forecasting systems.
1 Introduction
The forecasting community has seen a rapid growth of TSFMs (Kottapalli et al., 2025). These models are typically pre-trained on large collections of heterogeneous time series and then adapted to generate forecasts on unseen datasets (Ma et al., 2024; Oreshkin et al., 2021; Bommasani et al., 2022). So far, we have found approaches spanning a diverse set of architectures: non-Transformer models such as Tiny Time Mixers (Ekambaram et al., 2024); encoder-only designs including Moment (Goswami et al., 2024) and Moirai (Woo et al., 2024); decoder-only architectures such as Timer-XL (Liu et al., 2025a), Time-MOE (Shi et al., 2025), ToTo (Cohen et al., 2024), Timer (Liu et al., 2024a), TimesFM (Das et al., 2023), and Lag-Llama (Rasul et al., 2023); approaches that adapt pretrained LLMs to time series forecasting, including Chronos (Ansari et al., 2024), AutoTimes (Liu et al., 2024b), LLMTime (Gruver et al., 2024), Time-LLM (Jin et al., 2023), and FPT (Lu et al., 2021); and models deployed as a web API such as TimeGPT-1 (Garza et al., 2023a). Each line of work emphasizes different inductive biases and pretraining strategies, showing strong performance on benchmarks while highlighting trade-offs in scalability, interpretability, and robustness.
While this proliferation of models has expanded the options available to a seasoned forecaster when determining the best approach for a given use case, it has also increased complexity: since a particular laboratory or research group develops each model, each comes with its own API, training pipeline, and evaluation conventions. Beyond these differences, models also vary in their data input requirements (e.g., univariate vs. multivariate, patch vs. non-patch tokenization) and in their learning curve complexity, with some demanding extensive compute and expertise to deploy or use effectively. This fragmentation makes it difficult to compare models fairly, and even harder to integrate them seamlessly into production forecasting systems (Bergmeir, 2024). Recent large-scale evaluations such as GIFT-Eval (Aksu et al., 2024) highlight these challenges by showing in practice how different design choices across foundation models complicate reproducibility and cross-benchmark comparisons.
Meanwhile, the broader AI community has witnessed an explosion in the use of LLMs to orchestrate tasks across domains Brown et al. (2020); Vaswani et al. (2017). In particular, the agentic paradigm, where an LLM serves as a controller that plans, reasons, and executes tasks (Liu et al., 2024c), has been increasingly adopted in applications spanning text, code, image, video, and audio generation (Sapkota et al., 2025). Agents allow end-users to specify high-level goals in natural language, while the underlying system orchestrates specialized models and tools to achieve them.
Within time series, however, agentic approaches are still very new. For example, Cai et al. (2025) introduces a scalable benchmarking framework to evaluate AI agents in machine learning engineering tasks related to time series, such as data handling and code translation, rather than forecasting itself. In parallel, multimodal efforts such as Williams et al. (2025) highlight another emerging intersection between time series and LLMs: integrating numerical signals with natural language context to produce more reliable forecasts. Together, these efforts underscore both the promise and the nascency of agentic and LLM-augmented approaches in the forecasting domain. Despite these first steps, no prior work has unified TSFMs under an agentic interface.
In this paper, we introduce TimeCopilot [^1] TimeCopilot is available via pip install timecopilot and documentation at https://timecopilot.dev., the first open-source framework designed specifically for agentic forecasting, see Figure 1. TimeCopilot is a generative forecasting agent where LLMs act as reasoning engines to orchestrate multiple TSFMs, as well as classical statistical, machine learning, and deep learning methods, through a single unified API. This design eliminates fragmented dependencies across research groups and automates the full forecasting pipeline, from data preparation and model selection to ensembling and evaluation, while providing natural language explanations and supporting direct queries about the future. By bridging the agent paradigm with a unified hub of forecasting models, TimeCopilot provides a practical path toward reproducible, interactive, and accessible agentic forecasting workflows.
from timecopilot import TimeCopilot
tc = TimeCopilot(
llm="openai:gpt-4o"
result = tc.forecast(
df=df,
query="how many air passengers are expected in the next 12 months?",
print(result.output.user_query_response)
"The total expected air passengers for the next 12 months is approximately 5,919."
Figure 1: (Right) Overall TimeCopilot’s architecture. (Left) TimeCopilot Agent API usage.
2 Design Principles
TimeCopilot was designed to automate classical forecasting pipelines (Armstrong, 2001; Hyndman et al., 2024a). Traditionally, a forecaster explores a time series, extracts relevant information (features, plots, or diagnostics), formulates hypotheses about which models may perform well, and then uses cross-validation and benchmarking to select the final model (Makridakis et al., 2020). This cycle of exploration, hypothesis, and empirical validation has long been considered the cornerstone of forecasting practice.
TimeCopilot leverages LLMs in two ways: (i) to decide how to act at each step of this process, using their capacity to orchestrate multiple tools and reason over diverse signals, and (ii) to help forecasters explain both the decisions made during model selection and the forecasts themselves in natural language, making the results transparent and accessible. This combination of automation and explainability makes TimeCopilot the first agentic interface tailored to time series forecasting.
As Figure 1 shows, the framework is LLM-agnostic: it can interface with commercial APIs (e.g., OpenAI (Brown et al., 2020), Anthropic) as well as open-source models such as DeepSeek (DeepSeek-AI et al., 2025) or LLaMA (Touvron et al., 2023), giving users flexibility to choose models according to cost, availability, or deployment constraints.
2.1 TimeCopilot Agent
The TimeCopilot Agent orchestrates the forecasting workflow in three structured steps: (i) Time Series Feature Analysis: computes focused diagnostics (trend, seasonality, stationarity) that directly inform model choice (Garza et al., 2023b; Hyndman et al., 2024b) (ii) Model Selection and Evaluation: proposes candidate models starting from simple statistical baselines, documents their assumptions, evaluates them through cross-validation, and escalates to more complex models only when needed. (Hyndman et al., 2024a) (iii) Final Model Selection and Forecasting: chooses the best model based on metrics, generates forecasts while interpreting patterns, uncertainty, and reliability. Beyond automating these steps, the agent is designed for explainability: users can query not only the forecasts but also the reasoning behind each decision. This transparency differentiates TimeCopilot from black-box agentic systems and is critical for trust in forecasting applications.
2.2 TimeCopilot Forecaster
The TimeCopilot Forecaster executes the models proposed by the agent. It provides access to the largest unified hub of TSFMs under a single API  [^2] A complete and regularly updated list of models is available at https://timecopilot.dev/model_hub/, alongside statistical, machine learning, and deep learning methods. This hub removes the need for users to manage conflicting dependencies or learn multiple fragmented interfaces. The current set of supported TSFMs includes Chronos (Ansari et al., 2024), Moirai (Woo et al., 2024), Sundial (Liu et al., 2025b), TabPFN (Hoo et al., 2025), TiRex (Auer et al., 2025), TimeGPT (Garza et al., 2023a), TimesFM (Das et al., 2023), and Toto (Cohen et al., 2024).
Alongside foundation models, TimeCopilot integrates a wide suite of statistical baselines such as ADIDA, AutoARIMA, AutoETS, Theta, and SeasonalNaive (Garza et al., 2022), as well as Prophet (Taylor and Letham, 2018). For machine learning methods, TimeCopilot includes AutoLGBM (Nixtla contributors, 2023). For neural networks, it incorporates AutoNHITS (Challu et al., 2023) and AutoTFT (Lim et al., 2021) via the NeuralForecast library (Olivares et al., 2022).
Finally, TimeCopilot supports ensemble techniques, such as MedianEnsemble, which can combine forecasts from heterogeneous models (statistical, ML, and foundation) to improve robustness (Bates and Granger, 1969). This breadth enables direct comparison and flexible combinations, ensuring forecasters can benchmark and deploy the most suitable approach with minimal effort.
3 Benchmarks and API Usage Example
To illustrate the capabilities of TimeCopilot  we provide both benchmark results and minimal usage examples.
Figure 2: Performance of TimeCopilot and baseline models on the GIFT-Eval benchmark (Aksu et al., 2024). Lower values indicate better forecast performance. The top-15 models are shown, separated by non-leaking (top row) and all models (bottom row).
3.1 Benchmarks on GIFT-Eval
Figure 2 shows the performance of foundation models on the GIFT-Eval benchmark Aksu et al. (2024), which spans 24 datasets, 144k+ time series, and 177M data points across multiple domains and frequencies. An ensemble of foundation models, built with TimeCopilot, achieved the best probabilistic forecast performance (CRPS; Gneiting and Katzfuss 2014) and the second-best point forecast performance (MASE; Hyndman and Koehler 2006) among non-leaking models. Importantly, GIFT-Eval distinguishes between models that may have had test leakage and those trained strictly on unseen data [^3] Live results can be found at https://huggingface.co/spaces/Salesforce/GIFT-Eval ; TimeCopilot remained competitive against both categories while maintaining reproducibility and keeping computational costs low (about \$24 of GPU-distributed inference).
The TimeCopilot results rely on its MedianEnsemble feature  [^4] https://timecopilot.dev/api/models/ensembles. (Petropoulos and Svetunkov, 2020; Bates and Granger, 1969), used to combine three state-of-the-art foundation models (Moirai (Woo et al., 2024), Sundial (Liu et al., 2025b), and Toto (Cohen et al., 2024)) with isotonic regression (Barlow and Brunk, 1972) to ensure monotonic quantiles for probabilistic forecasting. This approach provides robustness against outliers and model-specific biases.  [^5] The full experiment can be reproduced at https://timecopilot.dev/experiments/gift-eval.
These results highlight three aspects of TimeCopilot: (i) It can orchestrate multiple foundation models through a unified interface. (ii) It provides state-of-the-art results in both probabilistic and point forecasting. (iii) It enables reproducible and cost-effective large-scale experimentation.
3.2 API Usage Example
TimeCopilot provides two main entry points: the Agent (end-to-end orchestration with natural language explanations and queries about the future) see Figure 1, and the Forecaster (direct control of specific models under a unified API), see Appendix 5 for more details. This simple interface allows users to switch seamlessly between agent-driven automation and manual benchmarking of specific models, while keeping the same data structure and evaluation utilities.
4 Conclusion and Future Work
We presented TimeCopilot, the first open-source agentic interface and framework that combines multiple TSFMs with LLMs to make the forecasting workflow explainable, transparent, and accessible to a wider audience. TimeCopilot addresses the challenge of unifying diverse forecasting models under a single API, while leveraging LLMs to automate pipeline decisions and provide natural language explanations for both model choices and forecasts.
TimeCopilot’s future work encompasses: i) Integrating with Model Context Protocol (MCP) enabling seamless interaction with external tools and services through standardized interfaces, ii) Expanding use cases beyond academic benchmarks to domains such as energy, climate, finance, and supply chain forecasting, and iii) Extending forecasting capabilities to support hierarchical forecasting, multivariate forecasting with multivariate explanations, and advanced evaluation settings such as coherence across multiple aggregation levels.
Acknowledgments
We would like to express our gratitude to Martin Juárez and Salomón Márquez for their helpful contributions and support.